OpenStudy (anonymous):

Where is the vertex of the parabola? y = x2 + 2x + 1 A. Above the x-axis B. Below the x-axis C. On the x-axis D. On the y-axis

OpenStudy (s3a):

y = x^2 + 2x + 1 y = (x + 1)^2 y = (x - h)^2 + k y = (x - -1)^2 + 0 Therefore, the vertex is at (h,k) = (-1,0) which means C is the correct answer.

OpenStudy (anonymous):

Wait what??? IS it A or C??????

OpenStudy (s3a):

vjordillo, you did + 1 + 1 instead of + 1 - 1 and x^2 + 2x + 1 = (x+1)^2

OpenStudy (anonymous):

SOOOOO IS IT A OR C?

OpenStudy (s3a):

It's C. I checked with software just now to see that I am right.